# Towards Efficient Training and Evaluation of Robust Models against $l_{0}$ Bounded Adversarial Perturbations

### v1

@@ Abstract
This work studies sparse adversarial perturbations bounded by $l_{0}$ norm. We propose a white-box PGD-like attack method named sparse-PGD to effectively and efficiently generate such perturbations. Furthermore, we combine sparse-PGD with a black-box attack to comprehensively and more reliably evaluate the models’ robustness against $l_{0}$ bounded adversarial perturbations. Moreover, the efficiency of sparse-PGD enables us to conduct adversarial training to build robust models against sparse perturbations. Extensive experiments demonstrate that our proposed attack algorithm exhibits strong performance in different scenarios. More importantly, compared with other robust models, our adversarially trained model demonstrates state-of-the-art robustness against various sparse attacks. Codes are available at https://github.com/CityU-MLO/sPGD.
@@ 1 Introduction
Deep learning has been developing tremendously fast in the last decade. However, it is shown vulnerable to adversarial attacks: imperceivable adversarial perturbations [1, 2] could change the prediction of a model without altering the input’s semantic content, which poses great challenges in safety-critical systems. Among different kinds of adversarial perturbations, the ones bounded by $l_{\infty}$ or $l_{2}$ norms are mostly well-studied [3, 4, 5] and benchmarked [6], because these adversarial budgets, i.e., the sets of all allowable perturbations, are convex, which facilitates theoretical analyses and algorithm design. By contrast, we study perturbations bounded by $l_{0}$ norm in this work. These perturbations are sparse and quite common in physical scenarios, including broken pixels in LED screens to fool object detection models and adversarial stickers on road signs to make an auto-driving system fail [7, 8, 9, 10, 11].
@@
However, constructing $l_{0}$ bounded adversarial perturbations is challenging as the corresponding adversarial budget is non-convex. Therefore, gradient-based methods, such as projected gradient descent (PGD) [4], usually cannot efficiently obtain a strong adversarial perturbation. In this regard, existing methods to generate sparse perturbations [12, 13, 14, 15, 16] either cannot control the $l_{0}$ norm of perturbations or have prohibitively high computational complexity, which makes them inapplicable for adversarial training to obtain robust models against sparse perturbations. The perturbations bounded by $l_{1}$ norm are the closest scenario to sparse perturbations among convex adversarial budgets defined by an $l_{p}$ norm. Nevertheless, adversarial training in this case [17, 18] still suffers from issues such as slow convergence and instability. Jiang et al. [19] demonstrates that these issues arise from non-sparse perturbations bounded by $l_{1}$ norm. In other words, $l_{1}$ adversarial budget still cannot guarantee the sparsity of the perturbations. Thus, it is necessary to study the case of $l_{0}$ bounded perturbations.
@@
In this work, we propose a white-box attack named sparse-PGD (sPGD) to effectively and efficiently generate sparse perturbations bounded by $l_{0}$ norm. Specifically, we decompose the sparse perturbation $\bm{\delta}$ as the product of a magnitude tensor ${\bm{p}}$ and a binary sparse mask ${\bm{m}}$: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ , where ${\bm{p}}$ and ${\bm{m}}$ determine the magnitudes and the locations of perturbed features, respectively. We adopt PGD-like algorithms to update ${\bm{p}}$ and ${\bm{m}}$. However, it is challenging to directly optimize the binary mask ${\bm{m}}$ in the discrete space. We thereby introduce an alternative continuous variable $\widetilde{{\bm{m}}}$ to approximate ${\bm{m}}$ and update $\widetilde{{\bm{m}}}$ by gradient-based methods, $\widetilde{{\bm{m}}}$ is then transformed to ${\bm{m}}$ by projection to the discrete space. Due to the sparsity of ${\bm{m}}$ by the projection operator, the gradient of ${\bm{p}}$ is sparse, which may lead to slow convergence by coordinate descent. Therefore, we can remove the projection operator in the backpropagation to obtain the unprojected gradient of ${\bm{p}}$. We use both the original sparse gradient and the unprojected gradient of ${\bm{p}}$ to boost the attack performance. Moreover, we design a random reinitialization mechanism to enhance the exploration capability for the mask ${\bm{m}}$. On top of sPGD, we propose sparse-AutoAttack (sAA), which is the ensemble of the white-box sPGD and another black-box sparse attack, for a more comprehensive and reliable evaluation against $l_{0}$ bounded perturbations. Through extensive experiments, we show that our method exhibits better performance than other attacks.
@@
    We propose an effective and efficient white-box attack algorithm named sparse-PGD (sPGD) to generate $l_{0}$ bounded adversarial perturbations.
@@
    sPGD achieves the best performance among white-box sparse attacks. We then combine it with a black-box sparse attack to construct sparse-AutoAttack (sAA) for more comprehensive robustness evaluation against $l_{0}$ bounded adversarial perturbations.
@@
2 Preliminaries
We use image classification as an example, although the proposed methods are applicable to any classification model. Under $l_{p}$ bounded perturbations, the robust learning aims to solve the following min-max optimization problem.
@@ 3 Related Works
Sparse Attacks: For perturbations bounded by $l_{0}$ norm, directly adopting vanilla PGD as in Eq. (2) leads to suboptimal performance due to the non-convexity nature of the adversarial budget: $\mathrm{PGD_{0}}$ [13], which updates the perturbation by gradient ascent and project it back to the adversarial budget, turns out very likely to trap in the local maxima. Different from $\mathrm{PGD_{0}}$, CW L0 [35] projects the perturbation onto the feasible set based on the absolute product of gradient and perturbation and adopts a mechanism similar to CW L2 [35] to update the perturbation. SparseFool [12] and GreedyFool [15] also generate sparse perturbations, but they do not strictly restrict the $l_{0}$ norm of perturbations. If we project their generated perturbations to the desired $l_{0}$ ball, their performance will drastically drop. Sparse Adversarial and Interpretable Attack Framework (SAIF) [36] is similar to our method in that SAIF also decomposes the $l_{0}$ perturbation into a magnitude tensor and sparsity mask, but it uses the Frank-Wolfe algorithm [37] to separately update them. SAIF turns out to get trapped in local minima and shows poor performance on adversarially trained models. Besides white-box attacks, there are black-box attacks to generate sparse adversarial perturbations, including CornerSearch [13] and Sparse-RS [16]. However, these black-box attacks usually require thousands of queries to find an adversarial example, making it difficult to scale up to large datasets.
@@
4 Methods
In this section, we introduce sparse-PGD (sPGD): a white-box attack that generates sparse perturbations. Similar to AutoAttack [34, 18], we further combine sPGD with a black-box attack to construct sparse-AutoAttack (sAA) for more comprehensive and reliable robustness evaluation.
In the end, we incorporate sPGD into the framework of adversarial training to boost the model’s robustness against sparse perturbations.
@@ 4.1 Sparse-PGD (sPGD)
$\max_{\|\bm{\delta}\|_{0}\leq k,0\leq{\bm{x}}+\bm{\delta}\leq 1}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})=\max_{{\bm{p}}\in{\mathcal{S}}_{\bm{p}},{\bm{m}}\in{\mathcal{S}}_{\bm{m}}}\mathcal{L}(\bm{\theta},{\bm{x}}+{\bm{p}}\odot{\bm{m}}).$ (3)
The feasible sets for ${\bm{p}}$ and ${\bm{m}}$ are ${\mathcal{S}}_{\bm{p}}=\{{\bm{p}}\in\mathbb{R}^{h\times w\times c}|0\leq{\bm{x}}+{\bm{p}}\leq 1\}$ and ${\mathcal{S}}_{\bm{m}}=\{{\bm{m}}\in\{0,1\}^{h\times w\times 1}|\|{\bm{m}}\|_{0}\leq k\}$, respectively. Similar to PGD, sPGD iteratively updates ${\bm{p}}$ and ${\bm{m}}$ until finding a successful adversarial example or reaching the maximum iteration number.
@@
Update Sparsity Mask ${\bm{m}}$: The sparsity mask ${\bm{m}}$ is binary and constrained by its $l_{0}$ norm. Directly optimizing the discrete variable ${\bm{m}}$ is challenging, so we update its continuous alternative $\widetilde{{\bm{m}}}\in\mathbb{R}^{h\times w\times 1}$ and project $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{\bm{m}}$ to obtain ${\bm{m}}$ before multiplying it with the magnitude tensor ${\bm{p}}$ to obtain the sparse perturbation $\bm{\delta}$. Specifically, $\widetilde{{\bm{m}}}$ is updated by gradient ascent. Projecting $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{{\bm{m}}}$ is to set the $k$-largest elements in $\widetilde{{\bm{m}}}$ to $1$ and the rest to 00. In addition, we adopt the sigmoid function to normalize the elements of $\widetilde{{\bm{m}}}$ before projection.
@@
$\widetilde{{\bm{m}}}$ $\longleftarrow\widetilde{{\bm{m}}}+\beta\cdot\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}/||\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}||_{2},$ (5)
@@
Backward Function: Based on Eq. (3), we can calculate the gradient of the magnitude tensor ${\bm{p}}$ by $\nabla_{\bm{p}}{\mathcal{L}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot{\bm{m}}$ and use $g_{{\bm{p}}}$ to represent this gradient for notation simplicity. At most, $k$ non-zero elements are in the mask ${\bm{m}}$, so $g_{\bm{p}}$ is sparse and has at most $k$ non-zero elements. That is to say, we update at most $k$ elements of the magnitude tensor ${\bm{p}}$ based on the gradient $g_{\bm{p}}$. Like coordinate descent, this may result in suboptimal performance since most elements of ${\bm{p}}$ are unchanged in each iterative update. To tackle this problem, we discard the projection to the binary set ${\mathcal{S}}_{{\bm{m}}}$ when calculating the gradient and use the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ to update ${\bm{p}}$. Based on Eq. (6), we have $\widetilde{g}_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot\sigma(\widetilde{{\bm{m}}})$. The idea of the unprojected gradient is inspired by training pruned neural networks and lottery ticket hypothesis [55, 56, 57, 58]. All these methods train importance scores to prune the model parameters but update the importance scores based on the whole network instead of the pruned sub-network to prevent the sparse update, which leads to suboptimal performance.
@@
Algorithm 1 Sparse-PGD
1:  Input: Clean image: ${\bm{x}}\in[0,1]^{h\times w\times c}$; Model parameters: ${\bm{\theta}}$; Max iteration number: $T$; Tolerance: $t$; $l_{0}$ budget: $k$; Step size: $\alpha$, $\beta$; Small constant: $\gamma=2\times 10^{-8}$
@@
3:  for $i=0,1,...,T-1$ do
4:     ${\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$
@@
13:     ${\bm{d}}=g_{\widetilde{{\bm{m}}}}/(||g_{\widetilde{{\bm{m}}}}||_{2})$ if $||g_{\widetilde{{\bm{m}}}}||_{2}\geq\gamma$ else 00
@@
15:     if attack succeeds then
16:        break
17:     end if
18:     if $||\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))-{\bm{m}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
19:        Random initialize $\widetilde{{\bm{m}}}$
20:     end if
21:  end for
22:  Output: Perturbation: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$

### v3

# Sparse-PGD: A Unified Framework for Sparse Adversarial Perturbations Generation

@@ Abstract
This work studies sparse adversarial perturbations, including both unstructured and structured ones. We propose a framework based on a white-box PGD-like attack method named Sparse-PGD to effectively and efficiently generate such perturbations. Furthermore, we combine Sparse-PGD with a black-box attack to comprehensively and more reliably evaluate the models’ robustness against unstructured and structured sparse adversarial perturbations. Moreover, the efficiency of Sparse-PGD enables us to conduct adversarial training to build robust models against various sparse perturbations. Extensive experiments demonstrate that our proposed attack algorithm exhibits strong performance in different scenarios. More importantly, compared with other robust models, our adversarially trained model demonstrates state-of-the-art robustness against various sparse attacks.
@@ 1 Introduction
Deep learning has been developing tremendously fast in the last decade. However, it is shown vulnerable to adversarial attacks: imperceivable adversarial perturbations [1, 2] could change the prediction of a model without altering the input’s semantic content, which poses great challenges in safety-critical systems. Among different kinds of adversarial perturbations, the ones bounded by $l_{\infty}$ or $l_{2}$ norms are mostly well-studied [3, 4, 5] and benchmarked [6], because these adversarial budgets, i.e., the sets of all allowable perturbations, are convex, which facilitates theoretical analyses and algorithm design. By contrast, we study sparse perturbations in this work, including both unstructured and structured ones which are bounded by $l_{0}$ norm and group $l_{0}$ norm, respectively. These perturbations are quite common in physical scenarios, including broken pixels in LED screens to fool object detection models and adversarial stickers on road signs to make an auto-driving system fail [7, 8, 9, 10, 11].
@@
However, constructing such adversarial perturbations is challenging as the corresponding adversarial budget is non-convex. Therefore, gradient-based methods, such as projected gradient descent (PGD) [4], usually cannot efficiently obtain a strong adversarial perturbation. In this regard, existing methods to generate sparse perturbations [12, 13, 14, 15, 16, 17] either cannot control the $l_{0}$ norm or the group $l_{0}$ norm of perturbations or have prohibitively high computational complexity, which makes them inapplicable for adversarial training to obtain robust models against sparse perturbations. The perturbations bounded by $l_{1}$ norm are the closest scenario to $l_{0}$ bounded perturbations among convex adversarial budgets defined by an $l_{p}$ norm. Nevertheless, adversarial training in this case [18, 19] still suffers from issues such as slow convergence and instability. Jiang et al. [20] demonstrates that these issues arise from non-sparse perturbations bounded by $l_{1}$ norm. In other words, $l_{1}$ adversarial budget guarantees still cannot the sparsity of the perturbations. Furthermore, existing works investigating structured sparse perturbations [21, 22, 23, 16, 17] only support generating a single adversarial patch, lacking flexibility and generality. Thus, it is necessary but challenging to develop a unified framework for both unstructured and structured sparse adversarial perturbations.
@@
In this work, we propose a white-box attack named Sparse-PGD (sPGD) to effectively and efficiently generate unstructured and structured sparse perturbations. For unstructured sparse perturbations, we decompose the perturbation $\bm{\delta}$ as the product of a magnitude tensor ${\bm{p}}$ and a binary sparse mask ${\bm{m}}$: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$ , where ${\bm{p}}$ and ${\bm{m}}$ determine the magnitudes and the locations of perturbed features, respectively. Although ${\bm{p}}$ can be updated by PGD-like methods, it is challenging to directly optimize the binary mask ${\bm{m}}$ in the discrete space. We thereby introduce an alternative continuous variable $\widetilde{{\bm{m}}}$ to approximate ${\bm{m}}$ and update $\widetilde{{\bm{m}}}$ by gradient-based methods, $\widetilde{{\bm{m}}}$ is then transformed to ${\bm{m}}$ by projection to the discrete space. To further boost the performance, we propose the unprojected gradient of ${\bm{p}}$ and random reinitialization mechanism. For structured sparse perturbations, we study group $l_{0}$ norm and its approximated version based on the group norm proposed in [24]. The structured sparse perturbation bounded by approximated group $l_{0}$ norm can be thus decomposed as ${\bm{\delta}}={\bm{p}}\odot{\bm{m}}={\bm{p}}\odot\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}}),1\right)$, where the binary group mask ${\bm{v}}$ determines the positions of the groups to be perturbed and the customized binary kernel ${\bm{k}}$ determines the pattern of groups. That is to say, we decompose the structured sparse perturbations into their positions and the patterns. Furthermore, we leverage transposed convolution ($\mathrm{TConv}$) and clipping operations to map the positions of the groups ${\bm{v}}$ to the positions of perturbed features ${\bm{m}}$. Similar to the unstructured cases, we introduce the continuous $\widetilde{{\bm{v}}}$ and update it by gradient-based methods. Ultimately, we manage to transform the problem of optimizing the structured sparse mask ${\bm{m}}$ into the problem of optimizing the group mask ${\bm{v}}$ bounded by $l_{0}$ norm, which can be resolved in the framework of sPGD. The whole pipeline is illustrated in Figure 2. On top of sPGD, we propose Sparse-AutoAttack (sAA), which is the ensemble of the white-box sPGD and another black-box sparse attack, for a more comprehensive and reliable evaluation against both unstructured and structured sparse perturbations. Through extensive experiments, we show that our method exhibits better performance than other attacks.
@@
    We propose an effective and efficient white-box attack algorithm named Sparse-PGD (sPGD) which can be utilized to generate both unstructured and structured sparseadversarial perturbations.
@@
    sPGD achieves the best performance among white-box sparse attacks. We then combine it with a black-box sparse attack to construct Sparse-AutoAttack (sAA) for more comprehensive robustness evaluation against sparse adversarial perturbations.
@@
Preliminaries: We use image classification as an example, although the proposed methods are applicable to any classification model. Under $l_{p}$ bounded perturbations, the robust learning aims to solve the following min-max optimization problem.
@@ 2 Related Works
Sparse Attacks: For perturbations bounded by $l_{0}$ norm, directly adopting vanilla PGD as in Eq. (2) leads to suboptimal performance due to the non-convexity nature of the adversarial budget: $\mathrm{PGD_{0}}$ [13], which updates the perturbation by gradient ascent and project it back to the adversarial budget, turns out very likely to trap in the local maxima. Different from $\mathrm{PGD_{0}}$, CW L0 [40] projects the perturbation onto the feasible set based on the absolute product of gradient and perturbation and adopts a mechanism similar to CW L2 [40] to update the perturbation. SparseFool [12] and GreedyFool [15] also generate sparse perturbations, but they do not strictly restrict the $l_{0}$ norm of perturbations. If we project their generated perturbations to the desired $l_{0}$ ball, their performance will drastically drop. Sparse Adversarial and Interpretable Attack Framework (SAIF) [41] is similar to our method in that SAIF also decomposes the $l_{0}$ perturbation into a magnitude tensor and sparsity mask, but it uses the Frank-Wolfe algorithm [42] to separately update them. SAIF turns out to get trapped in local minima and shows poor performance on adversarially trained models. Besides white-box attacks, there are black-box attacks to generate sparse adversarial perturbations, including CornerSearch [13] and Sparse-RS [17]. However, these black-box attacks usually require thousands of queries to find an adversarial example, making it difficult to scale up to large datasets. In addition to the unstructured adversarial perturbations mentioned above, there are several works discussing structured sparse perturbations, including universal adversarial patch for all data [21, 22] and some image-specific patches [23, 16, 17]. However, these works only support generating a single adversarial patch, which lacks of flexibility and generality.
@@
3 Unstructured Sparse Adversarial Attack
In this section, we introduce Sparse-PGD (sPGD) for generating unstructured sparse perturbations. Its extension that generates structured sparse perturbations is introduced in Sec. 4. Similar to AutoAttack [39, 19], we further combine sPGD with a black-box attack to construct sparse-AutoAttack (sAA) for more comprehensive and reliable robustness evaluation.
@@ 3.1 Sparse-PGD (sPGD)
$\max_{\|\bm{\delta}\|_{0}\leq\epsilon,0\leq{\bm{x}}+\bm{\delta}\leq 1}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})=\max_{{\bm{p}}\in{\mathcal{S}}_{\bm{p}},{\bm{m}}\in{\mathcal{S}}_{\bm{m}}}\mathcal{L}(\bm{\theta},{\bm{x}}+{\bm{p}}\odot{\bm{m}}).$ (3)
The feasible sets for ${\bm{p}}$ and ${\bm{m}}$ are ${\mathcal{S}}_{\bm{p}}=\{{\bm{p}}\in\mathbb{R}^{h\times w\times c}|0\leq{\bm{x}}+{\bm{p}}\leq 1\}$ and ${\mathcal{S}}_{\bm{m}}=\{{\bm{m}}\in\{0,1\}^{h\times w\times 1}|\|{\bm{m}}\|_{0}\leq\epsilon\}$, respectively. Similar to PGD, sPGD iteratively updates ${\bm{p}}$ and ${\bm{m}}$ until finding a successful adversarial example or reaching the maximum iteration number.
@@
Update Sparsity Mask ${\bm{m}}$: The sparsity mask ${\bm{m}}$ is binary and constrained by its $l_{0}$ norm. Directly optimizing the discrete variable ${\bm{m}}$ is challenging, so we update its continuous alternative $\widetilde{{\bm{m}}}\in\mathbb{R}^{h\times w\times 1}$ and project $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{\bm{m}}$ to obtain ${\bm{m}}$ before multiplying it with the magnitude tensor ${\bm{p}}$ to obtain the sparse perturbation $\bm{\delta}$. Specifically, $\widetilde{{\bm{m}}}$ is updated by gradient ascent. Projecting $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{{\bm{m}}}$ is to set the $\epsilon$-largest elements in $\widetilde{{\bm{m}}}$ to $1$ and the rest to 00. In addition, we adopt the sigmoid function to normalize the elements of $\widetilde{{\bm{m}}}$ before projection.
@@
$\widetilde{{\bm{m}}}$ $\longleftarrow\widetilde{{\bm{m}}}+\beta\cdot\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}/\|\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}\|_{2},$ (5)
@@
Backward Function: Based on Eq. (3), we can calculate the gradient of the magnitude tensor ${\bm{p}}$ by $\nabla_{\bm{p}}{\mathcal{L}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot{\bm{m}}$ and use $g_{{\bm{p}}}$ to represent this gradient for notation simplicity. At most, $\epsilon$ non-zero elements are in the mask ${\bm{m}}$, so $g_{\bm{p}}$ is sparse and has at most $\epsilon$ non-zero elements. That is to say, we update at most $\epsilon$ elements of the magnitude tensor ${\bm{p}}$ based on the gradient $g_{\bm{p}}$. Like coordinate descent, this may result in suboptimal performance since most elements of ${\bm{p}}$ are unchanged in each iterative update. To tackle this problem, we discard the projection to the binary set ${\mathcal{S}}_{{\bm{m}}}$ when calculating the gradient and use the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ to update ${\bm{p}}$. Based on Eq. (6), we have $\widetilde{g}_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot\sigma(\widetilde{{\bm{m}}})$. The idea of the unprojected gradient is inspired by training pruned neural networks and lottery ticket hypothesis [60, 61, 62, 63]. All these methods train importance scores to prune the model parameters but update the importance scores based on the whole network instead of the pruned sub-network to prevent the sparse update, which leads to suboptimal performance.
@@
Algorithm 1 Sparse-PGD for $l_{0}$ Bounded Perturbations
1:  Input: Clean image: ${\bm{x}}\in[0,1]^{h\times w\times c}$; Model parameters: ${\bm{\theta}}$; Max iteration number: $T$; Tolerance: $t$; $l_{0}$ budget: $\epsilon$; Step size: $\alpha$, $\beta$; Small constant: $\gamma=2\times 10^{-8}$
@@
3:  ${\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$
4:  for $i=0,1,...,T-1$ do
@@
13:     ${\bm{d}}=g_{\widetilde{{\bm{m}}}}/\|g_{\widetilde{{\bm{m}}}}\|_{2}$ if $||g_{\widetilde{{\bm{m}}}}||_{2}\geq\gamma$ else 00
@@
15:     ${\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$
16:     if attack succeeds then
17:        break
18:     end if
19:     if $||{\bm{m}}-{\bm{m}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
20:        Random initialize $\widetilde{{\bm{m}}}$
21:     end if
22:  end for
23:  Output: Perturbation: $\bm{\delta}={\bm{p}}\odot{\bm{m}}$